OpenStudy (anonymous):

The volume of a cone is 300 cubic kilometers. If the radius of the base is 7 kilometers, what is the height of the cone? Use 3.14 for pi. Show your work and round to the nearest tenth.

OpenStudy (anonymous):

is this correct? Volume = (radius^2 * PI * height) / 3 300 = (7^2 * 3.14 * height) / 3 900= 49 * 3.14 * height 900 = 153.86 * height (about) 5.4 kilometers = height